# Holomorphic Yang-Mills fields on $B$-branes

### v1

@@ 2.1.1. Neighborhood of the diagonal
Let $Y$ be an analytic complex compact connected manifold. We denote by $i:\Delta\hookrightarrow Y\times Y$ the embedding of the diagonal. As a closed subvariety, $\Delta$ is defined by an ideal $\mathscr{I}$ of $\hat{\mathscr{O}}:={\mathscr{O}}_{Y\times Y}.$ We will consider the following ringed spaces
$(Y,\,{\mathscr{O}}),\;\;Y^{(1)}=\Big{(}\Delta,\,{\mathscr{O}}_{Y^{(1)}}:=\big{(}\hat{\mathscr{O}}/{\mathscr{I}}^{2}\big{)}|_{\Delta}\Big{)},\;\;(Y\times Y,\,\hat{\mathscr{O}}).$
$Y^{(1)}$ is the first infinitesimal neighborhood of $\Delta$.
@@
Given $f\in{\mathscr{O}}$ and $l\in\hat{\mathscr{O}}$, the product
(2.3) $(l+{\mathscr{I}}^{2})\cdot f=l\,(f\circ p_{2})+{\mathscr{I}}^{2}$
defines a right ${\mathscr{O}}$-module structure on $\hat{\mathscr{O}}/{\mathscr{I}}^{2}$. More explicitly, $(l(x,\,y)+{\mathscr{I}}^{2})\cdot f=l(x,\,y)f(y)+{\mathscr{I}}^{2}.$ Analogously, $f\cdot(l+{\mathscr{I}}^{2})=(f\circ p_{1})l+{\mathscr{I}}^{2}$ gives to $\hat{\mathscr{O}}/{\mathscr{I}}^{2}$ a left ${\mathscr{O}}$-module structure. However, the restrictions to ${\mathscr{I}}/{\mathscr{I}}^{2}$ of these left and right ${\mathscr{O}}$-module structures are equivalent.
The isomorphism ${\mathscr{O}}\oplus\Omega^{1}\simeq{\mathscr{O}}_{Y^{(1)}}$. The cotangent sheaf $\Omega^{1}:=\Omega^{1}_{Y}$ can be identified with the pullback $j^{-1}({\mathscr{I}}/{\mathscr{I}}^{2})$ [8, p. 407]. This identification is defined by the correspondence
(2.4) $\xi:j^{-1}({\mathscr{I}}/{\mathscr{I}}^{2})\to\Omega^{1},\;\;\;g+{\mathscr{I}}^{2}\mapsto\big{(}d_{x}g\big{)}|_{Y},$
where $d_{x}g$ is the exterior derivative of $g$ with respect to the variables $x$; i.e. considering $g$ as function of the variables $x$ and keeping the variables $y$ constant.
If $l\in\hat{\mathscr{O}}$, the correspondence $l+{\mathscr{I}}^{2}\mapsto l\circ j\oplus(d_{x}l)|_{Y}$ defines an isomorphism of right ${\mathscr{O}}$-modules
[…]
When the abelian sheaf ${\mathscr{O}}\oplus\Omega^{1}$ is endowed with the left ${\mathscr{O}}$-action
(2.6) $f\cdot(h\oplus\alpha):=fh\oplus(hdf+f\alpha),$
where $\alpha\in\Omega^{1}$, then $m$ in (2.5) is an isomorphism of left ${\mathscr{O}}$-modules. We summarize those well-known results in the following proposition.
@@
With the notations above introduced:
  1. (1)
    The ${\mathscr{O}}$-modules $j^{-1}({\mathscr{I}}/{\mathscr{I}}^{2})$ and $\Omega^{1}$ are canonically isomorphic.
  2. (2)
    The correspondence (2.5) defines an isomorphism between the right ${\mathcal{O}}$-modules $j^{-1}(\hat{\mathscr{O}}/{\mathscr{I}}^{2})$ and ${\mathcal{O}}\oplus{\Omega}^{1}$.
  3. (3)
    Equipped ${\mathcal{O}}\oplus{\Omega}^{1}$ with the left ${\mathscr{O}}$-structure defined in (2.6), the correspondence (2.5) is also an isomorphism of left ${\mathscr{O}}$-modules.
The exact sequence of right $\hat{\mathscr{O}}$-modules
$0\to{\mathscr{I}}/{\mathscr{I}}^{2}\to\hat{\mathscr{O}}/{\mathscr{I}}^{2}\to\hat{\mathscr{O}}/{\mathscr{I}}\to 0$
gives rise, by means the functor $j^{-1}$, the exact sequence of right ${\mathscr{O}}$-modules
(2.7) $0\to j^{-1}({\mathscr{I}}/{\mathscr{I}}^{2})\to j^{-1}(\hat{\mathscr{O}}/{\mathscr{I}}^{2})\to j^{-1}(\hat{\mathscr{O}}/{\mathscr{I}})\to 0,$
or in other terms $0\to\Omega^{1}\to{\mathscr{O}}\oplus\Omega^{1}\to{\mathscr{O}}\to 0$
2.1.2. The first jet sheaf
We denote by $\pi_{a}=p_{a}\circ i:\Delta\to Y$, for $a=1,2$. One has the following morphism of sheaves rings over $\Delta$ (see Subsection 2.1.1)
$\pi_{a}^{-1}{\mathscr{O}}\longrightarrow{\mathscr{O}}_{Y^{(1)}}={\hat{\mathscr{O}}}/{\mathscr{I}^{2}},\;\;\;\;h\mapsto h\circ\pi_{a}+{\mathscr{I}}^{2}.$
In particular, one can consider ${\mathscr{O}}_{Y^{(1)}}$ as a right ${\pi_{2}^{-1}{\mathscr{O}}}$-module.
Given ${\mathscr{F}}$ a left ${\mathscr{O}}$-module on $Y,$ its inverse image by $\pi_{2}$ is left ${\mathscr{O}}_{Y^{(1)}}$-module
$\pi_{2}^{*}({\mathscr{F}})={\mathscr{O}}_{Y^{(1)}}\otimes_{\pi_{2}^{-1}{\mathscr{O}}}\pi_{2}^{-1}{\mathscr{F}}.$
And the first jet sheaf ${\mathscr{J}}^{1}({\mathscr{F}})$ of ${\mathscr{F}}$ is left ${\mathscr{O}}$-module defined by
(2.8) ${\mathscr{J}}^{1}({\mathscr{F}})={\pi_{1}}_{*}\pi_{2}^{*}({\mathscr{F}}).$
Since $\pi_{1}^{*}$ is the left adjoint of ${\pi_{1}}_{*}$, one has
(2.9) ${\rm Hom}\,_{{\mathscr{O}}}\big{(}{\mathscr{F}},\,{\mathscr{J}}^{1}({\mathscr{F}})\big{)}={\rm Hom}\,_{{\mathscr{O}}_{Y^{(1)}}}\big{(}\pi_{1}^{*}{\mathscr{F}},\,\pi_{2}^{*}{\mathscr{F}}\big{)}.$
If ${\mathscr{H}}$ is a left ${\mathscr{O}}_{Y^{(1)}}$-module, the ${\mathscr{O}}$-structure on ${\pi_{1}}_{*}{\mathscr{H}}$ is defined by $f\cdot s=(f\circ\pi_{1})s$, where $f\in{\mathscr{O}}$ and $s\in{\mathscr{H}}$. On the other hand, the ${\mathscr{O}}$ action on
$k^{*}{\mathscr{H}}={\mathscr{O}}\otimes_{k^{-1}{\mathscr{O}}_{\Delta}}k^{-1}{\mathscr{H}}$
is determined by $f\cdot(1\otimes s)=f\otimes s=1\otimes gs=1\otimes(f\circ\pi_{1})s$, where $f=g\circ k$. Hence, $k^{*}{\mathscr{H}}$ and ${\pi_{1}}_{*}{\mathscr{H}}$ are isomorphic ${\mathscr{O}}$-modules. Thus, as $\pi_{2}\circ k={\rm id},$
(2.10) ${\mathscr{J}}^{1}({\mathscr{F}})=k^{-1}({\mathscr{O}}_{Y^{(1)}})\otimes_{{\mathscr{O}}}{\mathscr{F}}=j^{-1}(\hat{\mathscr{O}}/{\mathscr{I}}^{2})\otimes_{{\mathscr{O}}}{\mathscr{F}}.$
By Proposition 1, ${\mathscr{J}}^{1}({\mathscr{F}})$ is the abelian sheaf ${\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}})$ endowed with the following left ${\mathscr{O}}$-module structure
(2.11) $f\ast(\sigma\oplus\beta)=f\sigma\oplus(f\beta+df\otimes\sigma).$
One has the morphism of abelian sheaves
$\eta:{\mathscr{F}}\to{\mathscr{J}}^{1}({\mathscr{F}})={\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}}),\;\;\sigma\mapsto\sigma\oplus 0.$
And from (2.11), it follows
(2.12) $f\eta(\sigma)=\eta(f\sigma)+df\otimes\sigma.$
Summarizing, one has the exact sequence of ${\mathscr{O}}$-modules
[…]
Thus, taking into account (2.1) and (2.9), a gauge field on ${\mathscr{F}}$ will be an element $\psi\in{\rm Hom}({\mathscr{F}},\,{\mathscr{J}}^{1}({\mathscr{F}}))$ right inverse of $\pi$. That is, we give the following definition.
Definition 1.
A holomorphic gauge field on the coherent sheaf ${\mathscr{F}}$ is a right inverse of the ${\mathscr{O}}$-module morphism $\pi:{\mathscr{J}}^{1}({\mathscr{F}})\to{\mathscr{F}}.$
We introduce a piece of notation that will be used later on. Let ${\sf Forg}$ denote the forgetful functor from the category of ${\mathscr{O}}$-modules to the category of abelian sheaves over $Y.$ We set
$\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}})={\sf Forg}\big{(}{\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}})\big{)}.$
In this way ${\mathscr{J}}^{1}({\mathscr{F}})$ is the abelian sheaf $\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}})$ endowed with the ${\mathscr{O}}$-module structure defined in (2.11).
2.1.3. The Atiyah class.
The exact sequence (2.13) defines an element $a({\mathscr{F}})$ of ${\rm Ext}^{1}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))$, called the Atiyah class of ${\mathscr{F}}$. The exact sequence (2.13) does not split, in general, in the category of ${\mathscr{O}}$-modules. Thus, there will be exist gauge fields on ${\mathscr{F}}$ iff the Atiyah class $a({\mathscr{F}})$ vanishes.
If there exists a right inverse $\psi$ of $\pi$; then $\pi(\psi-\eta)=0$; and thus $\nabla:=\psi-\eta$ takes its values in $\Omega^{1}({\mathscr{F}})$. Moreover, by (2.12), for $f\in{\mathscr{O}}$ and $\sigma\in{\mathscr{F}}$
(2.14) $\nabla(f\sigma)=f\psi(\sigma)-f\eta(\sigma)+df\otimes\sigma=f\nabla(\sigma)+df\otimes\sigma;$
that is, for $\nabla$ holds the Leibniz’s rule satisfied by the connections on vector bundles.
We denote by ${\mathcal{C}}$ the set consisting of all morphisms of ${\mathbb{C}}_{Y}$-modules $\nabla:{\mathscr{F}}\to\Omega^{1}({\mathscr{F}})$ satisfying (2.14). An element $\nabla\in{\mathcal{C}}$ determines the morphism of ${\mathscr{O}}$-modules
$\kappa:{\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}})\to\Omega^{1}({\mathscr{F}}),\;\;(\sigma,\beta)\mapsto\beta-\nabla(\sigma).$
$\kappa$ is a left inverse of $\iota$. Hence, $\kappa$ defines a splitting of the extension (2.13) and thus it determines a right inverse of $\pi$; i. e. a holomorphic gauge field $\psi$ on ${\mathscr{F}}$. We have proved the following proposition.
Proposition 2.
If $a({\mathscr{F}})=0$, then the map $\nabla\mapsto\nabla+\eta$ defines a bijective correspondence between ${\mathcal{C}}$ and the set of holomorphic gauge fields on the coherent sheaf ${\mathscr{F}}$.
Proposition 3.

### v2

@@ 2.1.1. The first jet sheaf
Let $X$ be an analytic complex compact connected manifold. Let $i:\Delta\hookrightarrow X\times X$ be the embedding of the diagonal. As a closed submanifold, $\Delta$ is defined by an ideal $\mathscr{I}$ of $\hat{\mathscr{O}}:={\mathscr{O}}_{X\times X}.$ The first infinitesimal neighborhood of $\Delta$ is the following ringed space
$X^{(1)}=\Big{(}\Delta,\,{\mathscr{O}}_{X^{(1)}}:=\big{(}\hat{\mathscr{O}}/{\mathscr{I}}^{2}\big{)}|_{\Delta}\Big{)}$
[…]
Given ${\mathscr{F}}$ a left ${\mathscr{O}}_{X}$-module, its inverse image by $\pi_{2}$ is the left ${\mathscr{O}}_{X^{(1)}}$-module
$\pi_{2}^{*}({\mathscr{F}})={\mathscr{O}}_{X^{(1)}}\otimes_{\pi_{2}^{-1}{{\mathscr{O}}_{X}}}\pi_{2}^{-1}{\mathscr{F}}.$
And the first jet sheaf ${\mathscr{J}}^{1}({\mathscr{F}})$ of ${\mathscr{F}}$ is the left ${\mathscr{O}}$-module defined by (see [21, p. 505] [25, Sect. 2.4])
(2.2) ${\mathscr{J}}^{1}({\mathscr{F}})={\pi_{1}}_{*}\pi_{2}^{*}({\mathscr{F}}).$
We set $\Omega^{1}({\mathscr{F}}):=\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{F}},$ where $\Omega^{1}_{X}$ is the sheaf of holomorphic $1$-forms on $X$. The first jet sheaf ${\mathscr{J}}^{1}({\mathscr{F}})$ is the abelian sheaf ${\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}})$ endowed with the following left ${\mathscr{O}}_{X}$-module structure
(2.3) $f\cdot(\sigma\oplus\beta)=f\sigma\oplus(f\beta+df\otimes\sigma).$
One has the Atiyah exact sequence of ${\mathscr{O}}_{X}$-modules
[…]
where $\pi$ is the projection morphism.
Since $\pi_{1}^{*}$ is the left adjoint of ${\pi_{1}}_{*}$, one has
(2.5) ${\rm Hom}\,_{{\mathscr{O}}_{X}}\big{(}{\mathscr{F}},\,{\mathscr{J}}^{1}({\mathscr{F}})\big{)}={\rm Hom}\,_{{\mathscr{O}}_{X^{(1)}}}\big{(}\pi_{1}^{*}{\mathscr{F}},\,\pi_{2}^{*}{\mathscr{F}}\big{)}.$
We introduce a piece of notation that will be used later on. Let ${\sf Forg}$ denote the forgetful functor from the category of ${\mathscr{O}}_{X}$-modules to the category of abelian sheaves over $X.$ We set
$\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}}):={\sf Forg}({\mathscr{F})}\oplus{\sf Forg}(\Omega^{1}({\mathscr{F}})).$
In this way ${\mathscr{J}}^{1}({\mathscr{F}})$ is the abelian sheaf $\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}})$ endowed with the ${\mathscr{O}}_{X}$-module structure defined in (2.3). In short, we will write $\,{\mathscr{J}}^{1}({\mathscr{F}})={\mathscr{F}}\,\tilde{\oplus}\,\Omega^{1}({\mathscr{F}}).$
One has the morphism of ${\mathbb{C}}_{X}$-modules
(2.6) $\eta:{\mathscr{F}}\to\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}})={\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}}),\;\;\sigma\mapsto\sigma\oplus 0.$
And from (2.3), it follows
(2.7) $f\eta(\sigma)=\eta(f\sigma)+df\otimes\sigma,$
where $f\in\mathscr{O}_{X}$ is a function of the structure sheaf of $X$ and $\sigma$ a holomorphic section of ${\mathscr{F}}.$
Given $\phi\in{\rm Hom}_{\mathscr{O}_{X}}({\mathscr{F}},\,{\mathscr{J}}^{1}({\mathscr{F}}))$ a right inverse of $\pi$; that is, such that $\pi\circ\phi={\rm id}$. Then $p\circ(\phi-\eta)=0,$ where $p:\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}})\to{\mathscr{F}}$ is the projection morphism. Thus $\phi-\eta$ factors uniquely through ${\rm Ker}(p)=\Omega^{1}({\mathscr{F}})$, defining the morphism $\nabla$ in the following diagram in the category of ${\mathbb{C}}_{X}$-modules.
[…]
From (2.7), together with the fact that $\phi$ is an ${\mathscr{O}}_{X}$-morphism, it follows that $\nabla$ satisfies the Leibniz rule $\nabla(f\sigma)=\partial f\otimes\sigma+f\nabla(\sigma).$ Therefore, one can give a new definition of holomorphic connection equivalent to the one given in [17].
Definition 1.
The holomorphic connections on the coherent ${\mathscr{O}}_{X}$-module ${\mathscr{F}}$ are the elements of the following set
(2.8) $\{\phi\in{\rm Hom}_{\mathscr{O}_{X}}({\mathscr{F}},\,{\mathscr{J}}^{1}({\mathscr{F}}))\;|\;\pi\circ\phi={\rm id}\}.$
That is, a holomorphic connection $\phi$ is a splitting of the corresponding Atiyah exact sequence. On the other hand, by (2.5), $\phi$ is an element of ${\rm Hom}_{{\mathscr{O}}_{X^{(1)}}}\big{(}\pi_{1}^{*}{\mathscr{F}},\,\pi_{2}^{*}{\mathscr{F}}\big{)}$ which is the identity on $X.$ Thus, we recover (2.1).
2.2. Gauge fields on a $B$-brane
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a $B$-brane on the complex manifold $X$; that is, ${\mathscr{F}}^{\bullet}$ is an object of the category $D^{b}(X)$, the bounded derived category of coherent sheaves over $X$. The corresponding first jet complex ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ is defined by
${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}):=R\pi_{1*}L\pi_{2}^{*}{\mathscr{F}}^{\bullet}\simeq{\mathscr{O}}_{X^{(1)}}\otimes^{L}{\mathscr{F}}^{\bullet}.$
As ${\mathscr{O}}_{X^{(1)}}$ is the locally free module ${\mathscr{O}}_{X}\oplus\Omega^{1}_{X}$, then ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ is the complex of abelian sheaves $\sf{Forg}({\mathscr{F}}^{\bullet})\,\oplus\,\sf{Forg}(\Omega^{1}({\mathscr{F}}^{\bullet})),$ with ${\mathscr{O}}_{X}$-structure given by
(2.9) $f\cdot(\sigma^{\bullet}\oplus\beta^{\bullet})=f\sigma^{\bullet}\oplus(df\otimes\sigma^{\bullet}+f\beta^{\bullet}).$
$\,\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}):=\big{(}{\sf Forg}({\mathscr{F}}^{\bullet})\oplus{\sf Forg}(\Omega^{1}({\mathscr{F}}^{\bullet})),\,\,\delta_{J}^{\bullet}\,\big{)},$
where $\delta_{J}^{\bullet}=\delta^{\bullet}\oplus(1\otimes\delta^{\bullet}).$ The complex $\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ can be equipped with an ${\mathscr{O}}_{X}$-module structure (2.9). As $\delta$ is the coboundary operator of a complex ${\mathscr{O}}_{X}$-modules, then $\delta_{J}(f\cdot(-))=f\cdot(\delta_{J}(-)).$ Thus, ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ is the complex $\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ endowed with this ${\mathscr{O}}_{X}$-module structure (2.9).
One has the exact sequence of complexes of ${\mathscr{O}}_{X}$-modules
[…]
The morphisms in the derived category $D^{b}(X)$ defined by the arrows in the sequence (2.10) will also be denoted by $\iota$ and $\pi.$
According to Definition 1, we propose the following definition.
Definition 2.
A gauge field on ${\mathscr{F}}^{\bullet}$ is an element $\psi\in{\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))$, such that ${\pi\circ}\psi={\rm id}\in{\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,{\mathscr{F}}^{\bullet}).$
Remark 1.
The inclusion functor from the category $\mathbf{Coh}(X)$ of coherent sheaves on $X$ to its derived category $D^{b}(X)$ is fully faithful [9, p. 164]. Therefore, Definition 2 coincides with Definition 1 when it is applied to a brane consisting of only one nontrivial term ${\mathscr{F}}.$
Remark 2.
By the adjunction isomorphism
${\rm Hom}_{D^{b}(X^{(1)})}\big{(}L\pi_{1}^{*}{\mathscr{F}}^{\bullet},\,L\pi_{2}^{*}{\mathscr{F}}^{\bullet}\big{)}\simeq{\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))$
a gauge field on ${\mathscr{F}}^{\bullet}$ can be considered as an element of
${\rm Hom}_{D^{b}(X^{(1)})}\big{(}L\pi_{1}^{*}{\mathscr{F}}^{\bullet},\,L\pi_{2}^{*}{\mathscr{F}}^{\bullet}\big{)},$
which lifts the identity on $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$, in accordance with the observation at the beginning of Subsection 2.1 (see (2.1)).
Remark 3.
The gauge fields on ${\mathscr{F}}^{\bullet}$ are elements ${\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}));$ i.e., the are particular open strings between ${\mathscr{F}}^{\bullet}$ and ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ with ghost number 00 [2, Sect. 5.2], [18].
The exact sequence (2.10) determines a distinguished triangle
[…]
in the category $D^{b}(X)$ [16, p. 46], [9, p. 157]. As ${\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,.\,)$ is a cohomological functor, it follows that
@@ 2.2. Gauge fields on a $B$-brane
is an exact sequence.
Since $\mu(\psi)=\pi\circ\psi$, the existence of a holomorphic gauge field $\psi$ on ${\mathscr{F}}^{\bullet}$, is equivalent to $1_{{\mathscr{F}}^{\bullet}}\in{\rm Im}(\mu)={\rm Ker}(\nu)$. If $\psi$ and $\psi_{1}$ are gauge fields on the coherent sheaf ${\mathscr{F}^{\bullet}}$, then $\mu(\psi_{1}-\psi)=0$; i. e. $\psi_{1}-\psi\in{\rm Im}(\lambda)$. Hence, the set of holomorphic gauge fields on ${\mathscr{F}}^{\bullet}$, if nonempty, is an affine space with vector space ${\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\Omega^{1}({\mathscr{F}}^{\bullet})).$ Thus, the space of holomorphic gauge fields on ${\mathscr{F}}^{\bullet}$, is a finite dimensional affine space.
The Atiyah class $a({\mathscr{F}}^{\bullet})$ of ${\mathscr{F}}^{\bullet}$ is the image of $1\in{\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,{\mathscr{F}}^{\bullet})$ in ${\rm Ext}^{1}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))$. Hence, we have the following proposition.